The length of a rectangle is 6 inches more than three times the width. The perimeter is 84 inches
Answer by sachi(548)

The length of a rectangle is 6 inches more than three times the width
let width=x
then length=3x+6
The perimeter is 84 inches
so 2(x+3x+6)=84
or 4x+6=84/2=42
or 4x=42-6=36
or x=36/4=9
so the width= 9 in & length=3*9+6=33 in